OpenStudy (anonymous):

Radius of convergence for sin(2x)?

OpenStudy (anonymous):

not sure what you mean sin(2x) converges everywhere

OpenStudy (anonymous):

Wait, how do you know that?

OpenStudy (anonymous):

By using the Ratio Test?

OpenStudy (anonymous):

you mean the power series?

OpenStudy (anonymous):

the expansion is found by taking the expansion for \(\sin(x)\) and replacing \(x\) by \(2x\) radius of convergence is infinite

OpenStudy (anonymous):

Wait, could you go step by step to find the radius of convergence?

OpenStudy (anonymous):

do you know the expansion for \(\sin(x)\)?

OpenStudy (anonymous):

Yessir! series from n=1 to infinity of (-1)^(n+1) * x^(2n-1) / (2n-1)!

OpenStudy (anonymous):

ok then if you replace \(x\) by \(2x\) you get something like \[\sin(2x)=\sum_{n=0}^\infty{(-1)^n\frac{2^{2n+1}x^{2n+1}}{(2n+1)!}}\]

OpenStudy (anonymous):

use the ratio test and you will get it

OpenStudy (anonymous):

mine is different than yours only because i started at 0and you at 1, they are the same

OpenStudy (anonymous):

if you do the algebra for the ratio test you will get \[\frac{x^2}{2}\lim_{n\rightarrow \infty}{\frac{1}{(n+1)(2n+3)}}.\] which is pretty clearly 0 for all x

OpenStudy (anonymous):

I see, thank you !